### v2

@@ 3.2.3 An Asymptotic Test of RUM and its Generalizations
Statistical testing of random utility models is due to Kitamura and Stoye (2018), with important computational improvement by Smeulders et al. (2020). It has seen application to observational data (Deb et al., 2022) as well as lab experiments (Aguiar et al., 2022). Because we will slightly adapt the approach to our setting, and also for self-contained exposition, we next briefly explain it. We also remind the reader that RUM is here meant in a somewhat loose sense: It is not essential that the population is modelled as mixing a finite number of “admissible” types encoded in the columns of $\bm{A}$, not that admissibility coincides with standard economic rationality; hence, we can use the machinery to test nonstandard and, in particular, nested models.
@@
The null hypothesis
@@
$H_{0}:$ $\bm{BA}\nu=\pi,\exists\nu\geq 0$
$\Longleftrightarrow$ $\pi\in\mathcal{C}\equiv\{\bm{BA}\nu:\nu\geq 0\}$
$\Longleftrightarrow$ $\min_{\nu\geq 0}\{(\pi-\bm{BA}\nu)^{\prime}\Omega(\pi-\bm{BA}\nu)\}=0,$
where $\Omega$ is an arbitrary positive definite (and in practice diagonal) weighting matrix. Here, the first step is true because a vector $\nu$ can fulfil $\bm{BA}\nu=\pi$ only if its elements sum to $1$, thus we can relax the constraint on $\nu$. The last expression states that the residuals from projecting $\pi$ onto the cone $\mathcal{C}$ of rationalizable probabilities must equal 00 and suggests the scaled norm of the corresponding sample residuals as natural test statistic. Noting the similarity to specification tests in multiple equation models (Sargan, 1958; Hansen, 1982), call this statistic
@@
where $\hat{\pi}$ stacks estimated choice probabilities $(\hat{p}_{d}(A),1-\hat{p}_{d}(A))$ in an order corresponding to $\pi$ and $n$ is sample size.
The asymptotic distribution of $J_{n}$ is delicate to estimate because it depends discontinuously on where on $\mathcal{C}$ the true $\pi$ is. Valid inference therefore relies on a bootstrap procedure that not only recenters the empirical distribution to be on $H_{0}$ (as in the previous test) but that also shrinks $H_{0}$ in a locally unfavorable direction. Heuristically, the idea is once again similar to Generalized Moment Selection in moment inequalities, but the details differ because, while $\mathcal{C}$ is a finite polytope, we can only characterize it in terms of its vertices (the columns of $\bm{A}$) as opposed to its faces (which would correspond to linear inequality constraints and allow for Generalized Moment Selection if we had them).
Operationally, we recenter the bootstrap distribution of data at the projection onto the tightened cone
$\mathcal{C}_{\tau_{N}}\equiv\{\bm{BA}\nu:\nu\geq\bm{1}\cdot\tau_{n}/H\},$
where $\tau_{n}$ is a tuning parameter that we set equal to $\sqrt{\log(\underline{n})/\underline{n}}$ with $\underline{n}$ being expected sample cell size for any but the universal choice problem; $\bm{1}$ is a vector of $1$’s. [^16] Simple algebra reveals that $\underline{n}=\tfrac{2qn}{k(k+1)}$, where $k$ is the number of nondefault items in $X$ and $q$ is the number of choice problems other than the grand problem faced by each subject. Recall also that $H$ is the length of $\nu$, thus division by $H$ ensures that the above constraint is scaled by $\tau_{n}$ and not by the testing problem’s complexity. The essential feature here is that $\tau_{n}$ vanishes but slowly compared to estimation uncertainty in $\hat{\pi}$. Then the bootstrap analog of $J_{n}$ is defined as
@@
$\hat{\eta}_{\tau_{n}}$$\equiv$ $\arg\min_{\nu\geq\bm{1}\cdot\tau_{n}/H}\{(\hat{\pi}-\bm{BA}\nu)^{\prime}\Omega(\hat{\pi}-\bm{BA}\nu)\}$ (3.5)
and $\hat{\pi}^{*}$ is a simple nonparametric (clustered, as explained earlier) bootstrap analog of $\hat{\pi}$.
Validity of the resulting test under reasonable assumptions is established in Kitamura and Stoye (2018). We go beyond a completely straightforward implementation of their test because we do not weight questions equally. This possibility is anticipated in Kitamura and Stoye (2018) because they require only a diagonal weighting matrix $\Omega$ but has not, to our knowledge, been implemented before. It is motivated by the fact that, in our data, choice probabilities pertaining to the universal choice set $X$ will be estimated from a much larger sample cell, and therefore much more precisely, than others. Specifically, both the expected and also realized average relative cell size of the last versus other questions are determined by the experimental design, and it is this number that we will use, i.e. the weighting matrix will be
$\Omega=\left(\begin{array}[]{rrrrr}1&\cdots&0&0&0\\ \vdots&\ddots&\vdots&\vdots&\vdots\\ 0&\cdots&1&0&0\\ 0&\cdots&0&w&0\\ 0&\cdots&0&0&w\end{array}\right),$
where $w$ is easy to compute and takes a value of $w\approx 9$ in our experimental design. [^17] Indeed, $w=\tfrac{k(k+1)}{2q}$, with $(k,q)$ as in the preceding footnote. We do not weight questions by realized sample cell sizes, and we also do not estimate cell-specific variances by the binomial variance formula, in order to avoid data dependent weighting. In our application, these modifications would have minimal effect.
We close by observing that in general, this test can be expensive to compute. The experimental design that we settled on, partly to ensure reasonable sample cell sizes, is small enough so that this is not the case. However, in preparation, we also implemented an adaptation of the computational improvements in Smeulders et al. (2020). For the benefit of future users, these details are laid out in Appendix B.
@@ 4 An Application to Experimental Data
We next describe an application of the above methods to an experimental data set.
@@ 4.1 Experimental Design
The aim of the experiment is to collect data of the type needed to implement tests for the Min and RUM bounds. This means we require repeated observations of choices from a grand set of alternatives and a number of subsets, with all sets containing a default option. Moreover, based on the findings of Chernev et al. (2015), we want the choice problems to be non-trivial to increase the probability of finding choice overload. To this end, we ask subjects to make choices between amounts of experimental points expressed as sums, as in Caplin et al. (2011).
@@
Specifically, each non-default option is expressed as a sum of four numbers between zero and ten written in text. The value of choosing an option in experimental points is the value of the sum, and one experimental point is worth 50 cents. There are 12 non-default options in the grand set $X$. Following Caplin et al. (2011), we generated these as follows: First, we drew the value of an alternative was drawn from an exponential distribution truncated at 10 points with $\lambda=0.25$. Next, individual terms of sums were chosen stochastically ensuring neither the first nor the maximal summand was correlated with the total value of the option.
@@
Meanwhile, a default option was available in each choice set. This option provided 7 points and was the only option expressed as a single number. The default option also always appeared at the top of the screen, was pre-selected (i.e. the subject had to actively click on another option if they wanted to choose it) and the simplest alternative (as measured by the number of summands in the expression). Thus it fulfilled many criteria that have been proposed to define the “default” in choice overload situations (see for example Iyengar and Kamenica (2010)). Of the non-default options, 9 yielded a number of points strictly lower than the default, 1 yielded the same number of points as the default, and 2 yielded strictly more points. Figure 1 shows an example choice screen with the default and one other option. Figure 2 shows an example choice screen for the grand set.
Figure 1: A choice problem comparing one lottery to the default.
Figure 2: A choice problem comparing the grand set to the default.
@@
The collection of choice sets used in the experiment consisted of the grand set (12 alternatives plus the default), and all 1 and 2 alternative subsets (along with the default), for a total of 78 smaller choice sets. Based on the prior literature, we initially believed that 3 item sets are unlikely to trigger choice overload, while 13 item sets are likely to do so, if such an effect is present. Each subject was presented with 9 randomly selected small sets and the grand set, with the order of choice questions and the order of non-default options in each choice question randomized. One question was randomly selected for payment. Subjects in addition received a \$1 participation fee. A complete list of the choice alternatives and choice sets can be seen in Appendix A.2.181818The experiment was approved by the Institutional Review Boards of Columbia and Cornell Universities.
Note that we have selected a choice environment that should generate a good deal of heterogeneity in item quality, but in which there should be little preference heterogeneity. Based on the discussion of Section 3.1 we would anticipate the Min bound to be effective in identifying choice overload, whle the RUM bound may not necessarily provide further advantages.
@@
Subject Recruitment. The experiment was run on Amazon’s Mechanical Turk (MT) platform, a digital marketplace for work. This platform was chosen to make it easy to get data from a large number of subjects, each of whom needed to answer only a limited number of questions. “Requesters” post Human Intelligence Tasks, or “HITs,” which are usually simple, repetitive jobs that typically pay small sums for each completed task. Workers on MT view descriptions of the HITs, decide which to accept, and complete those HITs over the internet. In our case, subjects who accepted the HIT followed a link to an external webpage, where they completed the experiment. Upon completion they were given a randomly generated code, which was used to pay them the appropriate amount given their choices in the experiment.
Instead of posting HITs and recruiting subjects directly on MTurk, we used an online platform called CloudResearch. CloudResearch posts our experiment as a HIT on MTurk and recruits the subjects. All subjects recruited through CloudResearch’s platform are first screened through CloudResearch’s Sentry data validation system. Sentry is a short (1 minute) pre-experiment survey meant to ensure subjects are attentive, engaged, and ready to participate in the experiment. People who fail Sentry are routed away from the experiment. These pre-study validation surveys have been shown to be effective in increasing data quality (Chandler et al., 2019; Litman and Robinson, 2020).
@@
We recruited 2000 subjects for the experiment. In addition to CloudResearch’s screening, we restricted recruitment to MTurk workers who had completed over 1000 HITs and had an approval rating of over 97%. Following the instructions subjects had to answer two quiz questions to ensure the understood the instruction. 1833 subjects passed the quiz and completed the experiment, but one of these subjects had a browser related error which made their data unusable, leaving us with 1832 subjects’ data.
The resulting data provides 1832 choices from the grand set. We have at least 185 observations of choices from each of the small sets, with variation in sample size due to the random selection of choice questions.
4.2 Data Overview
Figure 3 presents a histogram showing the distribution of choice sets by the probability with which the default option was chosen from sets other than $X$. The distribution is bimodal. The choice sets with low probability of default choice are precisely the ones in which the best option yields more points than the default. The probability of default choice in $X$ is indicated by the dotted vertical line. It falls somewhat above the default choice probability of the lower group, but below that of the higher group. Appendix A.2 lists the default choice probabilities for each choice set.
@@ 4.3 Analysis
We next ask whether the data we collected are suggestive of choice overload by conducting tests in roughly increasing order of presumed power.
First, we establish that existing approaches would be unlikely to identify choice overload in our data set. The empirical frequency of default choice in the grand set, $\hat{p}_{d}(X)$, was $22.3\%$. The analogous empirical frequency across all subsets was $70.7\%$ and therefore considerably higher. A simple comparison of means would not reveal choice overload, but also masks extreme heterogeneity across choice items. A random selection of a single smaller set would also be unlikely to reveal choice overload. $23$ of $78$ small choice sets had default choice frequency below $22.3\%$, only $15$ of which significantly so at the 5% level (using Fisher’s exact test on disjoint samples but without controlling for multiple comparisons). Thus, if a researcher were to randomly select a small choice set to compare to the grand set, they would only find evidence for choice overload 19% of the time.
In a next step, we use the finite sample test to ask whether any of those $23$ frequencies is significantly below the analogous frequency in the grand choice set, taking into account sampling uncertainty. Strikingly, the answer is yes: The p-value against the null hypothesis that the grand choice default frequency is lowest is .00003.00003.00003.00003, and $5$ choice items are significantly below it at the $5\%$-level, where the p-values and significance levels just reported reflect Bonferroni adjustment across $78$ tests. [^19] The choice sets alluded to are numbers $(9,11,20,32,74)$ in order of appearance in Table 2. Similarly, the asymptotic test of the Min bound yields a p-value that we could not distinguish from 00 in $B=10000$ Monte Carlo simulations. Hence we find strong evidence of choice overload using this approach.
We next apply the test explained in Section 3.2.3 to see if our data is consistent with RUM. At a Monte Carlo replication size of $B=10000$, the p-value against this null hypothesis is equals 00 as well. This extremely strong rejection comes with a caveat: The p-value against all but the grand choice set equals .005.005.005.005. Therefore, “the data are consistent with RUM except that default choice from $X$ is too frequent” is not an appropriate description of our findings. However, this leaves open the possibility that choice overload had an effect already in some of the smaller choice sets.
@@
and equals $1$ if, and only if, sample proportions are rationalizable in terms of rational matrix $\bm{A}$. This fraction is .866.866.866.866 for the RUM, i.e. model (i) above, increases to .877.877.877.877 for model (ii), and equals .915.915.915.915 for model (iii).
Third, one may worry that model (iii) is just not very restrictive, while maybe models (i) and (ii) are. This cannot be literally true because our test statistics are positive in all three tests, hence empirical choice frequencies are not rationalizable under any model. But it could certainly be “morally true,” notably if all possible data sets are “close” to the model. In order to shed some light on this, we conducted some analyses inspired by Bronars (1987), Selten (1991), and Beatty and Crawford (2011). Specifically, for models (i)-(iii) we calculated the expected mean square error (MSE) when matching data that were generated from the uniform distribution on $[0,1]^{79}$ (i.e. choice probabilities drawn uniformly at random). The resulting values are 0.25 for model (i), 0.23 for (ii) and 0.21 for (iii). These numbers are all close to each other, and far above the MSE when these models are applied to the data (which range from 0.03 to 0.01), showing that even our most permissive model places significant restrictions on the data.
Finally, the data can be used to illustrate the difference between the Min bound, $p_{d}^{min}(X)$, and the RUM-based bound, $p_{d}^{RUM}(X)$. The empirical choice frequencies do not imply a well-defined $p_{d}^{RUM}(X)$ because they are not stochastically rationalizable even if choice set $X$ is ignored. However, one can easily compute the vector $\hat{\eta}$ of choice probabilities that are closest to the empirical ones while being rationalizable; in fact, this computation is a by-product of the statistical tests. This allows us to compute a feasible analog of $p_{d}^{RUM}(X)$. Its value in our data is $9.8\%$. To compare apples to apples, we report that the same $\hat{\eta}$ implies a Min bound of $p_{d}^{min}(X)=11.4\%$. Using the full implications of RUM is potentially much more informative than just testing monotonicity.

### v3

@@ 3.2.3 An Asymptotic Test of RUM and its Generalizations
Statistical testing of Random Utility Models is due to Kitamura and Stoye (2018), with important computational improvement by Smeulders et al. (2021). It has seen application to observational data (Deb et al., 2023) as well as lab experiments (Aguiar et al., 2023). Importantly, the approach only requires that the population is modeled as mixing a finite number of “admissible” types encoded in the columns of $\bm{A}$, not that admissibility coincides with standard economic rationality. Hence, we can use the machinery to test nonstandard and, in particular, nested models.
@@
A main insight in Kitamura and Stoye (2018) is that the null hypothesis
@@
$H_{0}:\min_{\nu\geq 0}\{(\pi-\bm{BA}\nu)^{\prime}\Omega(\pi-\bm{BA}\nu)\}=0,$
where $\Omega$ is an arbitrary positive definite (and in practice diagonal) weighting matrix. That is, the residuals from projecting $\pi$ onto the cone $\mathcal{C}$ of rationalizable probabilities must equal 00. This suggests the scaled norm of the corresponding sample residuals as test statistic. Noting the similarity to specification tests in multiple equation models (Sargan, 1958; Hansen, 1982), call this statistic
@@
where $\hat{\pi}$ is the sample analog of $\pi$ and $n$ is sample size.
Despite the superficial similarity to well-established methods, the asymptotic distribution of $J_{n}$ is hard to estimate because it depends discontinuously on where on $\mathcal{C}$ the true $\pi$ is. However, one main contribution of Kitamura and Stoye (2018) is precisely to overcome this problem. Following them, we approximate the distribution of $J_{n}$ by the one of a modified bootstrap analog
@@
$\hat{\eta}_{\tau_{n}}$$\equiv$ $\arg\min_{\nu\geq\bm{1}\cdot\tau_{n}/H}\{(\hat{\pi}-\bm{BA}\nu)^{\prime}\Omega(\hat{\pi}-\bm{BA}\nu)\},$ (3.5)
where $\tau_{n}$ is a tuning parameter that we set in accordance with the literature, [^12] Specifically, $\tau_{n}=\sqrt{\log(\underline{n})/\underline{n}}$, where $\underline{n}=\tfrac{2qn}{k(k+1)}$ is expected sample cell size for any but the universal choice problem; here, $k$ is the number of nondefault items in $X$ and $q$ is the number of “small” choice problems faced by each subject. Recall also that $H$ is the length of $\nu$, thus division by $H$ ensures that the above constraint is scaled by $\tau_{n}$ and not by the testing problem’s complexity. $\bm{1}$ is a vector of $1$’s, and $\hat{\pi}^{*}$ is a simple nonparametric (clustered, as explained earlier) bootstrap analog of $\hat{\pi}$.
We go beyond a completely straightforward implementation of Kitamura and Stoye’s (2018) test because we do not weight questions equally. This possibility is anticipated by Kitamura and Stoye (2018), who only a diagonal weighting matrix $\Omega$, but has not, to our knowledge, been implemented before. We use it because, in our data, choice probabilities pertaining to the universal choice set $X$ will be estimated from a much larger sample cell than others. We take this into account by weighting estimated probabilities for different questions by the expected sample cell size for that question; in practice, that means to weight all small questions equally and to put weight $w\approx 9$ on choice frequencies corresponding to the grand set. [^13] More precisely, $w=\tfrac{k(k+1)}{2q}$, with $(k,q)$ as in the previous footnote. We do not weight questions by realized sample cell sizes, and we also do not estimate cell-specific variances by the binomial variance formula, in order to avoid data dependent weighting. In our empirical application, these modifications would have minimal effect.
In general, this test can be expensive to compute. The experimental design that we settled on, partly to ensure reasonable sample cell sizes, is small enough so that this concern does not arise. However, in preparation, we also implemented an adaptation of the computational improvements in Smeulders et al. (2021). These details are laid out in Supplementary Appendix B.
@@ 4.1 Experimental Design
Our testing strategy requires observing choices from a grand set of alternatives and a number of subsets. Based on the findings of Chernev et al. (2015), we want the choice problems to be non-trivial to increase the probability of finding choice overload. To this end, we ask subjects to make choices between amounts of experimental points expressed as sums, where each non-default option is expressed as sum of four numbers between 00 and $10$ written in text. The value of choosing an option in experimental points is the value of the sum, and one experimental point is worth $50$ cents. There are $12$ non-default options in the grand set $X$. We generate these by first drawing the value of an alternative from an exponential distribution truncated at 10 points with $\lambda=0.25$, then randomly selecting individual terms of the sum so that neither the first nor the maximal summand was correlated with the total value of the option. (All of this follows Caplin et al. (2011).)
@@
Each choice set also contained a default option providing $7$ experimental points, expressed as a single number. This option furthermore appeared at the top of the screen and was pre-selected, so it was an obvious default choice. Of the non-default options, $9$ yielded a number of points strictly lower than the default, $1$ yielded the same number of points, and $2$ yielded strictly more points. Figure LABEL:fig:1(a) shows an example choice screen with the default and one other option. Figure LABEL:fig:1(b) shows an example choice screen for the grand set.
@@
The collection of choice sets consisted of the grand set and all subsets containing 1 or 2 alternatives along with the default, for a total of 78 smaller choice sets. Based on the prior literature, we initially believed that 3-item sets are unlikely to trigger choice overload, while 13-item sets are likely to do so if such an effect is present to begin with. Each subject was presented with 9 randomly selected small sets and the grand set, with the order of choice questions and the order of non-default options in each choice question randomized. One question was randomly selected for payment. Subjects in addition received a \$1 participation fee. A complete list of choice alternatives and choice sets is in Supplementary Appendix A.2.141414The experiment was approved by the Institutional Review Boards of Columbia and Cornell Universities.
we note that this is a situation in which we would anticipate item heterogeneity to outweigh preference heterogeneity – assuming people like more money to less, there should not be any preference heterogeneity. This means that, as per our discussion in Section 4.2, we would expect the Min bound to perform well relative to the RUM bound.
@@
Subject Recruitment. The experiment was run on Amazon’s Mechanical Turk (MT) platform. This platform was chosen to easily collect data from a large number of subjects, each answering a small number of questions. “Requesters” post Human Intelligence Tasks (“HITs”) – usually simple jobs that pay small sums for each completed task. Workers on MT view descriptions of the HITs, decide which to accept, and complete those HITs over the internet. In order to improve subject attentiveness and reduce the probability of responses from bots, we prescreened subjects using the platform CloudResearch. This has been shown to be effective in increasing data quality (Chandler et al., 2019; Litman and Robinson, 2020).
@@
We recruited 2000 subjects in April and May of 2022. In addition to CloudResearch’s screening, we restricted recruitment to MTurk workers who had completed over 1000 HITs and had an approval rating of over 97%. 1833 subjects passed a comprehension quiz [^15] Screenshots of the quiz and instructions can be found in Supplemental Appendix A.1. and completed the experiment, but one of these subjects had a browser-related error which made their data unusable, leaving us with 1832 subjects’ data. In addition to 1832 choices from the grand set, we have at least 185 observations of choices from each of the small sets, with variation in sample size due to the random selection of choice questions. Supplementary Appendix A.2 lists default choice frequencies for each choice set.
@@ 4.2 Analysis
We test for choice overload in roughly increasing order of presumed test power. [^16] To showcase applicability of our methods to the data structure that we assumed, we ignore some features of our data at hand, namely that we observed precise choices and that we have several choices per subject, although we do consider the effect of the latter on sampling uncertainty. First, classic tests from the literature would be unlikely to find evidence of choice overload. The frequency of default choice in the grand set was $22\%$; the analogous frequency across all subsets was $71\%$. A simple comparison of means would, therefore, not reveal choice overload, and a random selection of a single smaller set would also be unlikely to do so – only 15 of 78 (19%) small sets have default choice significantly below the grand set’s.
Next, we use the finite sample test to ask whether any single default frequency is significantly below the one in the grand choice set, taking into account sampling uncertainty and multiple testing. Strikingly, the answer is yes: After Bonferroni adjustment, the p-value against the null hypothesis that the grand choice set default probability is lowest is .00003.00003.00003.00003, and $5$ choice items are significantly lower at the $5\%$-level. Similarly, the asymptotic test of the Min bound yields a p-value that we could not distinguish from 00 in $B=10000$ Monte Carlo simulations. Hence, we find strong evidence of choice overload using this approach.
We next test consistency with RUM. At a Monte Carlo replication size of $B=10000$, the p-value equals 00 as well. This strong rejection comes with a caveat: The p-value against all but the grand choice set equals .005.005.005.005. Therefore, “the data are consistent with RUM except that default choice from $X$ is too frequent” is not an appropriate description of our findings. However, this leaves open the possibility that, contrary to our own prior, choice overload had an effect in some of the smaller choice sets.
@@
and equals $1$ if, and only if, sample frequencies are rationalizable in terms of the behavior encoded in $\bm{A}$. This fraction is .866.866.866.866 for the RUM, increases to .877.877.877.877 for model (ii), and equals .915.915.915.915 for model (iii).
Third, one may worry that model (iii) is just not very restrictive, while maybe models (i) and (ii) are. This cannot be literally true because our test statistics are positive in all three tests, hence empirical choice frequencies do not conform to any model. But it could be “morally true,” notably if all possible data sets are close to the model. We investigated this through analyses inspired by Bronars (1987), Selten (1991), and Beatty and Crawford (2011). Essentially, these approaches propose comparing the test statistic observed in the data to that that observed in randomly generated pseudo data that matches the true data in some regards (for example, the choice sets from which choices are observed). Specifically, for models (i)-(iii) we calculated the expected mean square error (MSE) when matching data that were generated from the uniform distribution on $[0,1]^{79}$. The resulting values are 0.25 for model (i), 0.23 for (ii) and 0.21 for (iii). These numbers are all close to each other and far above the MSE when these models are applied to the data (which range from 0.03 to 0.01), showing that even our most permissive model places significant restrictions on the data.
Finally, the data can be used to illustrate the difference between the Min and RUM bounds. While empirical choice frequencies do not imply a well-defined $p_{d}^{RUM}(X)$, one can easily compute the vector $\hat{\eta}$ of choice probabilities that are closest to the empirical ones while being rationalizable; in fact, this computation is a by-product of the statistical tests. This allows us to compute a feasible analog of $p_{d}^{RUM}(X)$. Its value in our data is $9.8\%$. To compare apples to apples, we report that the same $\hat{\eta}$ implies a Min bound of $p_{d}^{min}(X)=11.4\%$. Once again, using the full implications of RUM is potentially much more informative than just testing monotonicity.